Sol Summary Report – December 6th

  
Crew 216 Sol Summary Report 06-DEC-19
Sol: 12
Summary Title: Final sol in sim
Author: Marc Levesque
Mission Status: Active
Sol Activity Summary: Today was the last sol in sim for our crew's mission at MDRS. Andrew, Rich, and Evgenia conducted the final EVA in the morning to return rock samples and capture some final film footage. The rest of the day was devoted to taking a food inventory, packing, and starting the clean up process. The crew will officially go out of sim this evening at 2100.
Look Ahead Plan: Pack up personal gear, move Crew Car and personal vehicles to front of Hab for loading, refill Hab water tanks, and clean all MDRS areas.
Anomalies in work: N/A.
Weather: Clear and cold overnight.
Crew Physical Status: Healthy
EVAs: None
Reports to be filed: Sol Summary Report, Commander Report, Operations Report, GreenHab Report, Journalist Report, Research Report, and Photos.
Support Requested: None 

Crew 216 Sol Summary 05Dec2019

[title Sol Summary – December 05th]

Crew 216 Sol Summary Report 05-DEC-19
Sol: 11
Summary Title: Winding down
Author: Marc Levesque
Mission Status: Active
Sol Activity Summary: The crew spent the day trying to complete as many tasks as possible on their projects in anticipation of completing their MDRS mission. These included finishing Crew 215’s materials project, capturing more footage for Evgenia’s film, and Michael processing his drone footage and tending to GreenHab and Observatory activities, while others tended to household and administrative tasks. Tomorrow will be the crew’s last full Sol of sim, and work will begin on station clean up to ease the workload on Saturday after the crew goes out of sim on Firday evening.
Look Ahead Plan: Complete personal projects, including one EVA for returning rock samples and capturing final film footage, and start station clean up.
Anomalies in work: N/A.
Weather: Rain and cold overnight, then clearing.
Crew Physical Status: Healthy
EVAs: None
Reports to be filed: Sol Summary Report, Commander Report, Operations Report, GreenHab Report, Journalist Report, Research Report, Astronomy Report, and Photos.
Support Requested: None

Crew 216 Sol Summary 04Dec2019

[title Sol Summary – December 04th]

Crew 216 Sol Summary Report 04-DEC-19
Sol: 10
Summary Title: Phobos Peak and International Space Station
Author: Marc Levesque
Mission Status: Active
Sol Activity Summary: The morning’s EVA to Phobos Peak captured some great film footage for Evgenia’s project and allowed Michael an opportunity to do the same for his drone project. The landscape around MDRS is different no matter which direction you go, and the view from Phobos was spectacular. The afternoon EVA was short, less than 100 meters, to set up the radio in the Crew Commander’s vehicle for an attempt to conduct an amateur radio contact with the International Space Station. Alas, no one was listening, but it was certainly worth a shot. As usual, GreenHab and Astronomy activities were ongoing.
Look Ahead Plan: Starting to wrap up projects with no EVAs.
Anomalies in work: N/A.
Weather: Clear and cold.
Crew Physical Status: Healthy
EVAs: To Phobos Peak and Crew Commander’s vehicle
Reports to be filed: Sol Summary Report, Commander Report, Operations Report, GreenHab Report, Journalist Report, Research Report, EVA Report, Astronomy Report, and Photos.
Support Requested: None

[end]

Sol Summary – December 3rd

Crew 216 Sol Summary Report 03-DEC-19
Sol: 9
Summary Title: Ore and mineral survey project
Author: Marc Levesque
Mission Status: Active
Sol Activity Summary: Two EVAs were conducted in support of a radioactive ore and mineral survey. In the morning, Mike, Evgenia, and Rich conducted a reconnaissance of the area around Robert’s Rock Garden to identity optimal sites for the evening EVA. The second EVA team of Rich, Andrew and Mike then undertook the survey using an ultraviolet light source in shortwave UV mode.
Look Ahead Plan: Two EVAs are planned. One is to continue filming for Evgenia’s project at Phobos Peak. The second, weather permitting, is to set up a radio repeater to attempt an amateur radio contact with the International Space Station that will be directly overhead the MDRS in the early afternoon.
Anomalies in work: N/A.
Weather: Clear and cold.
Crew Physical Status: Healthy
EVAs: To Robert’s Rock Garden (twice).
Reports to be filed: Sol Summary Report, Commander Report, Operations Report, GreenHab Report, Journalist Report, Research Report, EVA Report, Astronomy Report, and Photos.
Support Requested: None

Sol Summary – December 2nd

Crew 216 Sol Summary Report 02-DEC-19

Sol: 8

Summary Title: Communications and Filming Projects

Author: Marc Levesque

Mission Status: Active

Sol Activity Summary: Two EVAs were conducted in support of ongoing projects. In the morning Rich, Andrew, and Marc climbed to the summit of the North Ridge to confirm and map an accessible route and a location for a possible radio repeater site, both of which were accomplished. In the afternoon, Andrew, Michael, and Evgenia drove to the end of Quarry Road and then proceeded on foot for a short ways into Lithe Canyon for the purpose of capturing drone and documentary film footage.

Look Ahead Plan: Two EVAs are planned in support of Mike’s radioactive mineral project.

Anomalies in work: N/A.

Weather: Clear and cold.

Crew Physical Status: Healthy

EVAs: To North Ridge and Lithe Canyon.

Reports to be filed: Sol Summary Report, Commander Report, Operations Report, GreenHab Report, Journalist Report, Research Report, EVA Report, Astronomy Report, and Photos.
Support Requested: None

Sol Summary – December 1st

Crew 216 Sol Summary Report 01-DEC-19
Sol: 7
Summary Title: Roker Film Crew
Author: Marc Levesque

Mission Status: Active

Sol Activity Summary: The crew spent the day hosting a film crew from Al Roker Entertainment. They were here to film one episode of a six-part series that will focus on different aspects of science. Upon arrival, they first did some filming outside the MDRS before joining us inside for some interviews and filming the Hab, GreenHab, Observatory, Science Dome, and RAM. The afternoon was spent filming an EVA with the episode host and then interviewing some individual crew members.

Look Ahead Plan: Two EVAs are planned to continue work on crew projects.

Anomalies in work: N/A.

Weather: Clear and cold.

Crew Physical Status: Healthy

EVA: To Kissing Camel Ridge for the Roker film crew.

Reports to be filed: Sol Summary Report, Commander Report, Operations Report, GreenHab Report, Journalist Report, Research Report, EVA Report, Astronomy Report, and Photos.

Support Requested: None

Sol Summary – November 30th

Crew 216 Sol Summary Report 30-NOV-19
Sol: 6
Summary Title: Rest and Reconnaissance
Author: Marc Levesque

Mission Status: Active

Sol Activity Summary: The crew spent the morning waiting for the ground to dry sufficiently to allow an afternoon EVA. This provided an opportunity for crew rest and pursuing projects for filming, the GreenHab, and astronomy. With sunny weather drying out the landscape, the afternoon EVA was approved and undertaken with caution. While the main objective of the EVA was to reach the summit of the North Ridge to determine a suitable repeater site was not achieved because of ground conditions, a reconnaissance and assessment of route alternatives to reach the top of the North Ridge was accomplished.

Look Ahead Plan: A film crew from Al Roker Entertainment will be arriving around 0900 to spend the day filming the MDRS, the crew, and its activities.

Anomalies in work: N/A.

Weather: Clearing and cold.

Crew Physical Status: Healthy

EVA: N/A

Reports to be filed: Sol Summary Report, Commander Report, Operations Report, GreenHab Report, Journalist Report, Research Report, EVA Report, Astronomy Report, and Photos.

Support Requested: None

Sol Summary – November 29th

Crew 216 Sol Summary Report 29-NOV-19
Sol: 5
Summary Title: Catch up day
Author’s name: Marc Levesque
Mission Status: Active
Sol Activity Summary: The crew took advantage of the windy and wet weather that included snow to catch up on indoor activities and more importantly, rest. Filming continued for Evgenia’s project, Andrew worked on the gypsum project, Michael spent time developing a GreenHab checklist, while others caught up on graduate research papers, completed household chores, or napped.
Look Ahead Plan: An EVA will be requested for tomorrow afternoon after the ground has a chance to dry out to check out the potential north repeater site that was not reached on Wednesday. We will also tend to some household chores in anticipation of the Roker film crew arriving on Sunday.
Anomalies in work: N/A.
Weather: Cold and wet overnight.
Crew Physical Status: Healthy
EVA: N/A
Reports to be filed: Sol Summary Report, Commander Report, Operations Report, GreenHab Report, Journalist Report, Research Report, and Photos.
Support Requested: None

Sol Summary – November 27th

Crew 216 Sol Summary Report 27-NOV-19
Sol: 3
Summary Title: More work on communications, film, and drone projects
Author’s name: Marc Levesque
Mission Status: Active
Sol Activity Summary: We conducted two EVAs today to continue work on the communications, filming, and drone projects. In the morning, Mike, Andrew, Evgenia, and I hiked to two potential repeater sites, one of which was used before. We ran out of time to visit the third and plan to visit it next week, but Evgenia was able to continue with her filming project. In the afternoon, Rich, Michael, and Evgenia drove south to complete yesterday’s EVA that was truncated by communication issues. This also allowed Evgenia and Michael to capture more footage.
Look Ahead Plan: We will conduct one EVA to perform rover cleaning if weather permits.
Anomolies in work: N/A.
Weather: Cold and clear overnight.
Crew Physical Status: Healthy
EVA: Two EVAs conducted.
Reports to be filed: Sol Summary Report, Commander Report, Operations Report, Astronomy Report, GreenHab Report, Journalist Report, Research Report, EVA Report, and Photos.
Support Requested: Trash pick up in rear airlock

Sol Summary – November 27th

Crew 216 Sol Summary Report 27-NOV-19
Sol: 3
Summary Title: More work on communications, film, and drone projects
Author’s name: Marc Levesque
Mission Status: Active
Sol Activity Summary: We conducted two EVAs today to continue work on the communications, filming, and drone projects. In the morning, Mike, Andrew, Evgenia, and I hiked to two potential repeater sites, one of which was used before. We ran out of time to visit the third and plan to visit it next week, but Evgenia was able to continue with her filming project. In the afternoon, Rich, Michael, and Evgenia drove south to complete yesterday’s EVA that was truncated by communication issues. This also allowed Evgenia and Michael to capture more footage.
Look Ahead Plan: We will conduct one EVA to perform rover cleaning if weather permits.
Anomolies in work: N/A.
Weather: Cold and clear overnight.
Crew Physical Status: Healthy
EVA: Two EVAs conducted.
Reports to be filed: Sol Summary Report, Commander Report, Operations Report, Astronomy Report, GreenHab Report, Journalist Report, Research Report, EVA Report, and Photos.
Support Requested: Trash pick up in rear airlock

Copyright © The Mars Society. All rights reserved. | Main Site